### v2

# An Empirical Study of Vulnerability Handling Times in CPython

Abstract
The paper examines the handling times of software vulnerabilities in CPython, the reference implementation and interpreter for the today’s likely most popular programming language, Python. The background comes from the so-called vulnerability life cycle analysis, the literature on bug fixing times, and the recent research on security of Python software. Based on regression analysis, the associated vulnerability fixing times can be explained very well merely by knowing who have reported the vulnerabilities. Severity, proof-of-concept code, commits made to a version control system, comments posted on a bug tracker, and references to other sources do not explain the vulnerability fixing times. With these results, the paper contributes to the recent effort to better understand security of the Python ecosystem.
Index Terms:
Python, defect, software security, software vulnerability, vulnerability disclosure, bug fixing times, bug severity
I Introduction
Underneath the Python programming language are so-called virtual machines that compile Python code into byte code before execution. These machines are embedded into the language’s interpreters. While there are many interpreters, including examples such as Cython and Jython, the reference implementation, CPython, is the most popular one. Over the years, many vulnerabilities have also been reported for this reference implementation [15]. The present paper examines how long the handling of these have taken, what factors explain the handling times, and how well these can be predicted.
By handling it is meant that addressing of vulnerabilities requires many distinct software engineering work tasks. A vulnerability needs to be obviously fixed, but a given fix needs to be also integrated into releases, often including distinct release branches. In addition, the vulnerability requires coordination between multiple parties [14, 27], which in the open source context include particularly so-called downstream distributors, such as Linux distributions. Coordination is also required with the non-profit MITRE corporation to get a Common Vulnerabilities and Exposures (CVEs) identifier. Although the Python Software Foundation (PSF), who as an organization is behind CPython, is a CVE numbering authority and can thus allocate CVEs on its own, it may be that additional coordination is still required with some CVEs before they are published by MITRE and later on archived into the National Vulnerability Database (NVD) [18], the world’s foremost vulnerability database. For these reasons, the paper concentrates on two distinct timelines within a vulnerability’s overall handling time: (a) the time required to fix a given vulnerability and (b) the time required for a CVE for it to be published. Hereafter, the former is known as fixing time and the latter as CVE coordination time.
The questions examined and the paper’s topic in general are easy to motivate. According to benchmarks, Python is the most popular programming language today [7], and because CPython is the most popular interpreter for the language, the vulnerabilities affecting the interpreter affect large user and deployment bases. In addition, as pointed out in the opening Section II, the handling times proxy not only software engineering effort but also security risks. A further motivating point is that the paper’s topic has not been examined previously, despite a large reference literature base on bug and vulnerability handling times, including their fixing times. In this regard, the paper contributes to the existing and related work by patching a small but notable knowledge gap.
The paper’s remaining structure is simple. After the already noted Section II on related work, the dataset and methods for examining it are elaborated in Section III. Then, the empirical results are presented in Section IV. Finally, Section V summarizes the conclusions reached, pinpoints some limitations, and discusses the implications particularly for further work.
II Related Work
There are two rather large branches of related work. The first branch is sometimes known as a vulnerability life cycle analysis [9, 16]. Like normal, non-security bugs, vulnerabilities are discovered, reported, coordinated, fixed, and archived to databases, among other things. These and other events that occur during a vulnerability’s life cycle allow to formulate different longitudinal research questions and setups for these.
For instance, a time difference between a date of discovering a vulnerability and a date when a discoverer first contacted a vendor affected by the vulnerability allows to approximate communication delays and potential communication obstacles in vulnerability disclosure [26]. As elaborated in Section III-B, the contacting dates, often also known as vulnerability disclosure dates, are important also in the present work because they set operational reference points for both timelines considered.
The PSF has also a specific vulnerability disclosure policy. In essence, a discoverer should privately contact a security team at the PSF, who then determines whether the issue reported is really a vulnerability, and if so, handles the required coordination with the discoverer privately, contemplates whether the vulnerability can be publicly discussed in a bug tracker, and then fixes the vulnerability, integrates the fix developed to releases, and releases security advisories for the vulnerability [21]. Against this backdrop, the fixing times considered proxy particularly the software engineering work required. Although not perfectly, both the fixing times and the CVE coordination times proxy also security risks; a long delay imply more potential for exploitation of a given vulnerability.
Similar timelines have been widely used in previous work [26, 27, 28]. The patch development aspect and the software engineering tenet imply that the branch and the paper too further interlace with a closely related empirical research domain that has examined and predicted bug fixing times [1, 3, 29, 34]. While also this domain essentially operates with time differences, bug tracking systems have prompted also more convoluted questions between a bug’s state changes, including a question of which bugs get reopened in bug tracking systems [35]. In contrast, the vulnerability life cycle branch usually, either explicitly or implicitly, maintains that a vulnerability’s life cycle is more or less a linear process.
The second research branch originates from the Python programming language itself. In particular, a lot of work has been done in recent years to examine the security of software written in Python. In addition to vulnerability detection in Python code [32], the branch has operated with an ecosystem-wide scope, examining particularly the packages distributed in the Python’s PyPI repository. Here, on one hand, many Python packages have been observed to suffer from various software quality issues, including real and potential security flaws [4, 25]. On the other hand, it has also been observed and argued that a probability of picking a safe, non-vulnerable package from PyPI is still relatively high [20]. To some extent, this observation aligns with results from time series analysis; only a recent past has been observed to be relevant for predicting a probability that a Python package’s current version is vulnerable [24]. By and large, but not entirely [15], this largely empirical branch of research has overlooked a fact that over the years a lot of vulnerabilities have been reported also for the language’s core, CPython itself. Because this reference implementation is written in the classical C language, also the vulnerabilities are somewhat or even very different than vulnerabilities in packages written in Python.
III Data and Methods
III-A Data
Although there are many tools for extracting data from software repositories, including those written in Python [31], the relatively small amount of vulnerabilities for CPython allowed to construct the dataset manually. Another, more practical reason for the manual construction is that CPython recently moved from a custom tracker [22] to a more systematic one using the Open Source Vulnerability (OSV) format [23]. The latter is hosted on GitHub. For the purposes of this paper, the old tracker had a benefit in that it recorded many distinct vulnerability handling dates explicitly. For this reason, the data contains only the $n=93$ vulnerabilities that were fully present in the old tracker. A few vulnerabilities from 2023 had to be excluded from the old tracker because they were not fully recorded due to the associated migration to the new tracker. Finally, it should be stressed that not all of the vulnerabilities that have affected CPython have been about the interpreter per se. As CPython bundles other software components, their vulnerabilities affect also the interpreter and its security. The Expat library is the prime example in this regard.
III-B Variables
III-B1 Dependent Variables
Two separate dependent variables are used for the empirical analysis. The first is:
$(\textit{Fixing Time})_{i}=(\textit{Last Commit})_{it_{a}}-(\textit{Disclosure})_{it_{c}},$ (1)
where $i=1,\ldots,n$ refers to the $i$:th vulnerability, $(\textit{Last Commit})_{it_{a}}$ to the last commit in a version control system that was made to fix the $i$:th vulnerability at date $t_{a}$, regardless of a branch, and $(\textit{Disclosure})_{it_{c}}$ to a date $t_{c}$ at which the $i$:th vulnerability was first disclosed to the PSF’s security team or otherwise made known to the CPython’s developers according to the meta-data from the old vulnerability tracker. If multiple persons reported a same vulnerability, the officially recorded disclosure date is still used. It should be also remarked that the CPython’s developers may have identified and categorized multiple vulnerabilities identified with a single CVE. In any case, $(\textit{Fixing time})_{i}$ is a typical count data variable approximating particularly the software engineering effort required to fix vulnerabilities.
The second dependent variable is operationalized as:
$(\textit{CVE Coordination Time})_{i}$ $=(\textit{CVE Publication})_{it_{b}}$ (2)
$-(\textit{Disclosure})_{it_{c}},$
where $(\textit{CVE Publication})_{it_{b}}$ is a date $t_{b}$ at which a CVE was published for the $i$:th vulnerability according to the meta-data from the old tracker. Four remarks are necessary about this variable. First, the vulnerabilities, as identified as such by the CPython’s developers, lacking CVEs had to be obviously excluded. Second, in case two or more CVEs were allocated for a single vulnerability, as identified as a single vulnerability in the old tracker, the one with the earliest date was used. Third, a restriction $(\textit{CVE Coordination Time})_{i}\geq 0~{}\forall~{}i$ was imposed, meaning that those vulnerabilities were excluded that had CVEs allocated already before the associated disclosure dates. These cases have supposedly happened due to a given discoverer or other reporter having obtained a CVE from MITRE on his or her own, without first contacting the PST’s security team. Fourth, as the variable’s name indicates, it is taken to proxy coordination of CVE identifiers, but it is also important in terms of security because many companies allegedly only deploy security patches fixing vulnerabilities identified with CVEs. Much work has also been done to help companies and others with patching prioritization [13, 33].
III-B2 Independent Variables
the independent, explanatory variables are as follows:
  1. 1.
    REPORTER: a set of dummy variables identifying reporters of the CPython vulnerabilities. Note that a reporter may or may not be the same person who originally discovered a given vulnerability. Reporters were identified by their unique real names or pseudonyms. The rationale for this independent variable builds on many existing studies, which have hypothesized and observed that reporters’ and discoverers’ skills and characteristics, including their communication skills, affect bug and vulnerability handling times [3, 10, 17, 26, 34].
  2. 2.
    SEVERITY: severity of the vulnerabilities patched in CPython, as measured with a $[0,10]$ interval-scaled variable based on the the Common Vulnerability Scoring System (CVSS v. 3) information; the higher a value, the more severe a vulnerability. If a CVE was missing, a value zero was used, and if NVD lacked CVSS data for a CVE, the CVSS (v. 2) information from the old tracker was used. Although existing results have not often been confirmatory [4, 17, 26], the severity of bugs and vulnerabilities has been a classical variable used to model and predict the associated handling times [1, 3, 10, 34]. However, the hypothesized direction of an effect is not entirely clear. On one hand, as developers may prioritize severe vulnerabilities, their patching times should be shorter. On the other hand, particularly severe vulnerabilities may be time-consuming to analyze and fix. In any case, in practice the CVSS (v. 3) severity information was obtained from hyperlinks in the CPython’s old tracker pointing directly to the NVD.
  3. 3.
    POC: a dummy variable indicating whether a reporter or some other person had included a proof-of-concept (POC) code for demonstrating a given vulnerability. The recording only counts POCs that were explicitly, as visible code, posted to the initial bug report for a vulnerability. In other words, POCs referenced by hyperlinks to external sources were excluded. The variable’s rationale comes from existing research on bug fixing times; POCs, reproducible tests, stack traces, screenshots, and associated things have been observed to shorten bug (and vulnerability) fixing times [11, 26, 29]. The reason seems clear: the more there is robust information available, the easier and hence faster it is to fix a vulnerability or a non-security bug.
  4. 4.
    COMMITS: the number of commits to a version control system that were required to patch a given vulnerability in all branches. The rationale is straightforward: a large amount of commits and thus effort should lengthen the timelines. If many commits were required, a vulnerability may have been particularly complex, difficult to interpret and debug, or otherwise time-consuming to fix. Although existing results are not entirely definite [28], code churn has also been observed to lengthen bug fixing times [34].
  5. 5.
    REFERENCES: a number of online references to external sources, counted in terms of hyperlinks posted to a given bug report. Although internal hyperlinks pointing to a tracker’s other bug reports or other elements were excluded together duplicate hyperlinks, the variable counts also hyperlinks posted by bots, including those referencing commits. The manually posted hyperlinks typically point toward bug trackers and other tracking systems used by other projects, including operating system vendors, although references to Python online documentation, standards, mailing lists, blogs, security companies, and many other things are present too. A similar variable has been used also in previous studies [27]. The rationale is that the vulnerability handling timelines might be shorter for vulnerabilities that are widely discussed and popularized on different communication channels and platforms. Though, it may also be that a widely discussed vulnerability is controversial or otherwise problematic, perhaps indicating a longer handling time already due to the time required to discuss and collectively interpret it.
  6. 6.
    COMMENTS: by again following existing research [10, 34], the number of comments that were posted to an initial bug report for a vulnerability. The rationale is similar to REFERENCES. It can be additionally remarked that the actual content of discussions on bug reports does not seem to affect bug fixing times [19]. Therefore, merely including a variable measuring a volume of discussions seems justified.
Four remarks are in order about these independent variables. First, it should be mentioned that during the period observed CPython had also used two bug tracking system, an internal one and a one hosted on GitHub. The presence of two trackers complicates the operationalization of COMMENTS and REFERENCES. In terms of the former variable, all comments were accounted for vulnerabilities that were handled in the internal tracker, whereas in the GitHub case the operationalization only counted comments identified as such by GitHub, meaning that mentions, references, and other entries in GitHub’s parlance were excluded. Then, in terms of REFERENCES, in addition to the exclusion of internal references, also cross-references between the two trackers were excluded. Second, in case a bug report was not referenced in the old tracker, POC, REFERENCES, and COMMENTS were all scored with a value zero. Third, even though hypothesized effects are not entirely unambiguous, all variables are well-justified in terms of existing research. Fourth and last, only six variables were operationalized even though plenty more would be easy to derive and operationalize. The reason is the small sample size.
III-C Methods
Broadly speaking, the literature has operated with two methodological approaches: classification and regression analysis. The former approach typically splits bug and vulnerability handling times into two categories; “short” and “long” or “fast” and “slow”  [1, 3, 10, 17, 34]. This approach suffers from an obvious limitation that a threshold for a split is more or less arbitrary. The second approach treats a patching time as a continuous count data variable. Therefore, the typical methodological choices include ordinary least squares with variable transformations [26], Poisson regression, negative binomial regression [25], quantile regression [27], and Cox’s proportional hazards regression [28]. The regression approach is used in the present work. Two estimators are used: a standard ordinary least squares (OLS) regression with a $\ln(x+1)$ transformation for the two dependent variables and a so-called Huber’s $M$-estimator with the same transformation. The latter belongs to a family of robust regression methods, and, therefore, it is generally much less sensitive to outliers. Without delving into the statistical details, which are well-documented [8], the $M$-estimator is computed with the rlm function from the MASS library for the R language.
IV Results
The empirical analysis can be started by taking a brief look at the volume of vulnerabilities across time. Thus, Fig. 1 displays the annual vulnerability counts. As can be seen, the amounts of vulnerabilities reported have steadily increased from the mid-2000s. The mean is about six vulnerabilities per year. Many of the vulnerabilities have also been rather severe, as can be concluded from Fig. 2. About 13% of the vulnerabilities have CVSS (v. 3) base scores higher than eight. The median is close to six. The severity values are higher than what have been reported for software written in Python [24], although not substantially higher. The slight divergence is presumably explained by the C programming language.
Figure 1: Publication Years According to Disclosure Dates
Figure 2: Severity of the Vulnerabilities (CVSS)
The two dependent variables are illustrated in Fig. 3 by using the noted logarithm transformation. As can be seen, both the vulnerability fixing times and the CVE coordination times have been rather lengthy. In terms of the former, the mean is $119$ days and the median is as long as $267$ days. The fixing times seem also longer than in software written in Python. Although methodology is not directly comparable, previous studies have reported that the median to fix vulnerabilities in packages distributed in PyPI is about two months [4]. Then, regarding the CVE coordination times, the median is $157$ days. Although methodology is again different, shorter timelines have been reported previously also in this regard [27]. Having said that, nine vulnerabilities satisfied a condition $(\textit{CVE Coordination Time})_{i}<0$, meaning that CVEs were already allocated before the vulnerabilities were disclosed to the PSF’s security team or the CPython’s developers publicly.
Figure 3: Handling Times
TABLE I: $F$-tests for the OLS Regressions ($F$-values)
Fixing time CVE coordination time
REPORTER ${}^{**\phantom{*}}$ ${}^{\phantom{***}}$
SEVERITY ${}^{\phantom{***}}$ ${}^{\phantom{***}}$
POC ${}^{\phantom{***}}$ ${}^{\phantom{***}}$
COMMITS ${}^{\phantom{***}}$ ${}^{\phantom{***}}$
REFERENCES ${}^{\phantom{***}}$ ${}^{\phantom{***}}$
COMMENTS ${}^{\phantom{***}}$ ${}^{\phantom{***}}$
$n$     ${}^{\phantom{***}}$ 69
R2 ${}^{\phantom{***}}$ ${}^{\phantom{***}}$
       ${}^{***}~{}\textmd{for}~{}p<0.001$, ${}^{**}~{}\textmd{for}~{}p<0.01$, and ${}^{*}~{}\textmd{for}~{}p<0.05$
Figure 4: Actual and Fitted Values from the OLS Regression for Fixing Times
Turning to the regression analysis, Table I summarizes the two OLS regressions with $F$-tests. Unlike conventional $t$-tests for determining the statistical significance of regression coefficients, a $F$-test determines the statistical significance of whole variable groups, which is particularly important regarding the dummy variables for the reporters of the vulnerabilities. And, indeed, REPORTER is the only statistically significant variable in the fixing time regression but not in the CVE coordination time regression. Despite this general lack of statistical significance, model performance is very good in both models. The fixing time regression yields a coefficient of determination as high as $0.923$. The good performance is further illustrated in Fig. 4. However, the performance is almost entirely because of the reporters. When fitting a new model with only REPORTER, a value $R^{2}=0.909$ is obtained.
Basic graphical diagnostic plots indicate only modest concerns. On one hand, normality of the residuals is an issue; Shapiro-Wilk tests [30] reject the null hypotheses that the residuals came from a normally distributed population in both OLS models. Though, this commonplace normality violation is debated, and the consequences may not be fatal at all [12]. On the other hand, heteroskedasticity is not a particular concern, unlike what has been reported in previous studies with similar study settings [26]. According to two Breusch-Pagan tests [6], the null hypotheses of homoskedasticity of the residuals remain in force in both OLS models. Recomputing the models with the Huber’s $M$-estimator does not change the conclusion; only some reporter dummy variables are statistically significant according to $t$-tests. Finally, it should be mentioned that the models estimated are close to being ill-defined, meaning that there are almost as many variables as there are observations. To this end, a $R^{2}=0.256$ value is obtained for an OLS model including explicit dummy variables for only those reporters who had reported at least two vulnerabilities; all others are cascaded into a group of “others”. Although the value is much lower than in the previous models, it is still quite reasonable for such a tiny regression model. Further including the remaining variables does not improve performance; all of these other variables are statistically insignificant, and $R^{2}$ actually decreases to $2.149$.
All in all, it can be concluded that some characteristics of people who report CPython vulnerabilities largely explain the associated vulnerability fixing times. A qualitative analysis would be likely needed to deduce about a subsequent why-question. Based on the manual skimming of the associated bug reports while constructing the dataset, it does not seem plausible that communication and related things would explain the finding. Rather, it may be that some people just write better vulnerability reports, perhaps providing things that developers typically consider helpful, including reproducible tests, sample code, and even patches [5, 11]. Alternatively, when considering the bundling of third-party libraries, it may also be that the faster fixing times have been due to people who are associated with Linux distributions or other open source communities. In such a case, verification, upstream coordination, and associated things may have already been done elsewhere prior to handling a vulnerability at the CPython’s development infrastructure.
V Discussion
V-A Conclusion
This short paper examined vulnerability handling times in CPython. Two variables were specifically examined: vulnerability fixing times and CVE coordination times. The paper’s conclusions can be summarized with the following points:
  • CPython has seen a steady flow of new vulnerabilities, and the arrival rates have slightly increased from the 2000s. Many of the vulnerabilities are rather severe.
  • Both the vulnerability fixing times and the CVE coordination times have been lengthy. The medians are $267$ and $157$ days, respectively. In other words, it has taken on average about nine months to fix the vulnerabilities.
  • Based on regression analysis, only an identification of persons who have reported vulnerabilities is relevant statistically. In fact, merely including this identification data yields very good statistical performance. Contrary to many closely related previous studies, severity, proof-of-concept code, commits, comments posted to a bug tracker, and references to other sources explain neither the vulnerability fixing times nor the CVE coordination times.
In addition to these brief conclusive three points, some limitations should be acknowledged. After elaborating these, a couple of points follow about the potential for further research.
V-B Limitations
Regarding limitations, the obvious needs to be explicitly mentioned: as only a single case was analyzed, nothing can be deduced about vulnerability handling times in other projects. Rather than trying to seek generalizability, which already in the open source context is difficult, if not even impossible, it might be more reasonable to address the limitation by examining other interpreters for interpreted programming languages. Such an examination might, or should, also reveal whether vulnerabilities are similar or different across interpreters.
Also construct validity and robustness of the data might be slightly threatened. As was described in detail in Section III, the presence of two vulnerability tracking systems complicated the data collection process. In addition, many concessions had to be made when operationalizing the variables for the analysis. However, these problems are fairly typical to the research domain. Already the fundamental abstract question of what counts as a vulnerability is not straightforward; a single CVE may reference multiple vulnerabilities and the other way around, at least according to the CPython’s developers. It should be also acknowledged that manual data collection often involves a degree of subjectivity; in this regard, particularly the POC independent variable may be subjected to criticism.
V-C Further Work
The paper’s main result—the importance of reporters and their characteristics—would deserve a closer look in further research. Why the vulnerability fixing times can be predicted so well by merely identifying the reporters of the CPython vulnerabilities? Although a similar observation has been made also in previous studies, the answers to the question have been only tentative and without much theoretical contributions. If expertise of reporters, including with respect to providing reproducible information, is as important as often seen in the literature [11], it might be also possible to make practical advances by providing better instructions to people about good vulnerability reporting practices. At the moment, the PSF’s policy [21] does not say anything about reproducibility, POCs, and other similar things related to vulnerability handling.
Another good question for further research would be to examine a third delay metric: the time required to integrate the fixes to releases. According to the CPython’s old vulnerability tracker, it seems that such integration too has taken a rather long time. The many software engineering work activities related to integration [2] may offer an explanation. However, it may well also be that new CPython releases are not pushed forward merely to address vulnerabilities. That is, the vulnerability fixes may be postponed to match existing release engineering schedules and plans. Whatever the explanation may be, the integration timelines are relevant in terms of actual security risks because users typically only get vulnerability fixes through releases, meaning that they are often running vulnerable Python interpreters until a new release is made. In fact, one could further extend the examination toward observing delays that occur until third-party distributors have integrated the CPython releases into their package managers.